OpenStudy (anonymous):

Can you show me how to solve this example? 3x^2+8=0

OpenStudy (anonymous):

\[3x^2=-8\] if you are working with real numbers you are done, because a square cannot be negative. if you are working with complex numbers then \[x^2=\frac{-8}{3}\] \[x=\pm\sqrt{-\frac{8}{3}}\] \[x=\pm\sqrt{\frac{8}{3}}i\]